OpenStudy (anonymous):

Write an equation in slope-intercept form for the line that satisfies the following condition. slope 4, and passes through (2, 20)

OpenStudy (anonymous):

Slope-intercept form is expressed as y=mx+b where m is the slope and b is the y-intercept. So, you have y=4x+b. To find b, plug in the coordinates given for x and y. 20=4(2)+b 20=8+b subtract 8 from both sides 12=b Now put it all together: y=4x+12 There's your answer.

OpenStudy (anonymous):

Thank You.